OpenStudy (nicoleeshe):

I have 2 points on a parabola 7,-10/6 and 8, -30/6 plus I have the y value of the vertex = 1. I am supposed to find the x intercepts but am stumped

OpenStudy (saadsword):

do you have the equation?

OpenStudy (nicoleeshe):

no, that is the problem. I need to develop the equation from the information given

OpenStudy (saadsword):

one second this might take a while

OpenStudy (nicoleeshe):

you would be my hero

OpenStudy (saadsword):

I think it is something along the lines of: y=ax^2+bx+c -10/6=a(7)^2+b(7)+c and -30/6=a(8)^2+b(8)+c then maybe you can try to find a relationship between the two equations and make them equal to each other. As of right now, im still thinking

OpenStudy (nicoleeshe):

I have looked at it numerous ways. As a vertex form it becomes: -10/6=a(7-h)^2 + 1 and -30/6=a(8-h)^2 +1 so 2 equations with 2 unknowns but the solution is incredibly unwieldy. Is that perhaps the only solution?

OpenStudy (saadsword):

ill leave for now and try to work this out on my note book, it is easier that way. ill be back

OpenStudy (nicoleeshe):

thanks I will check back

OpenStudy (phi):

why the strange y values ? isn't -30/6 just -5 ? and -10/6 is -5/3

OpenStudy (nicoleeshe):

I don't know...is there a clue in there somewhere

OpenStudy (phi):

I don't see any clue. This is tough, because it is non-linear also, there are different answers I can see drawing two different parabolas